Spencer Company expects to sell 60,000 units next year. Variable production costs are $4 per unit, and variable selling costs are 10% of the selling price. Fixed expenses are $115,000 per year, and the company has set a target profit of $50,000. Based on this information, the unit selling price should be:

A) $7.00

B) $10.75

C) $7.50

D) $6.75
Answer #1

Answer = C -  $7.5

(Selling Price per unit - Variable Costs)* Units Sold - Fixed Cost = Profit.

Let Selling Price be X.

So, Variable selling Costs =0.10X

(X- 4 - 0.10X) * 60000 - $115000 =$50000

0.9 X - 4 = ($50000 + $115000)/60000

.0.9X - 4 = 2.75

X = (2.75 + 4) /0.9

=$ 7.5

Comment if you have any doubts. Please Upvote the answer if you find it useful. Thankyou.